A square wire loop 15.0 cm on each side carries a clockwise current of 9.00 A .

Find the magnitude of the magnetic field at its center due to the four 1.50 -mm wire segments at the midpoint of each side.

Express your answer with the appropriate units.
Answer #1

Solution :-

magnetic field due to one side of wire,

B = u0*iL / 4*pi*r^2

B = 4*pi*10-7*9*1.1*5×10-3 / 4*pi*(0.075)2

B = 2.4*10-7 T

magnetic field at its center due to the four wire segments,

Btotal = 4B

Btotal = 4*(2.4*10-7)

Btotal = 9.6*10-7 T

Direction is into the page.